OpenStudy (anonymous):

If y varies directly as x and y=24 when x=8, then when y=15 x is

OpenStudy (anonymous):

since y is directly proportional to x, we have y=kx where k is a constant of proportionality.now when y=24,x=8 we have 24=k*8 which gives k=3 when y=15 x=y/k = 15/3 = 5

OpenStudy (anonymous):

is 5 the anwer than